In a study of the infestation of a particular species of lobster by two types of barnacles, the body lengths (in millimeters) of 10 randomly selected lobsters are measured.

78    67    66    63    62    61    58    58    52    51

Find a 95% confidence interval for the mean body length of this species of lobster. (Round your answers to three decimal places.)

____ mm to _____ mm
Answer #1

From sample data,

= 61.6 , S = 7.8202

95% Confidence Interval
X̅ ± t(α/2, n-1) S/√(n)
t(α/2, n-1) = t(0.05 /2, 10- 1 ) = 2.262
61.6 ± t(0.05/2, 10 -1) * 7.8202/√(10)
Lower Limit = 61.6 - t(0.05/2, 10 -1) 7.8202/√(10)
Lower Limit = 56.006
Upper Limit = 61.6 + t(0.05/2, 10 -1) 7.8202/√(10)
Upper Limit = 67.194

95% CI is from 56.006 to 67.194